# Sensing for Free: Learn to Localize More Sources than Antennas without Pilots

### v1

@@ IV-B Architecture of the Proposed Snap-TF
Next, we introduce the proposed Snap-TF architecture for multi-source DOA estimation, whose pseudo-codes are shown in Algorithm 1. We discuss the details as follows.
@@ IV-B1 Signals as Tokens
We initialize by treating each raw snapshot as an input token to the transformer network. Specifically, for an $M$-element SLA, each snapshot $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$ is split into its real and imaginary parts and concatenated into a real-valued $2M$-dimensional vector. Stacking $T$ such vectors forms the initial feature set $\mathbf{S}^{(0)}\in\mathbb{R}^{2M\times T}$, whose $t$-th column corresponds to the $t$-th snapshot. We bypass the tokenization procedures in vanilla transformers [45] to let the network directly process higher-order interactions between raw signal snapshots to learn the higher-order statistics that are important for DOA estimation under non-Gaussian modulated symbols. Also, we deliberately drop the positional encoding and masking commonly employed in vanilla transformers [45] to preserve the permutation invariance of the architecture with respect to the ordering of signal snapshots $\mathbf{y_{\bm{\Omega}}}(t)$. Since the order of snapshots should not affect the output, we treat $\mathbf{S}^{(0)}$ as the initial feature set whose columns are the set elements.
@@ IV-B2 Self-Attention Blocks for Higher-Order Interactions
Snap-TF employs a stack of transformer-inspired single-head self-attention blocks to capture the higher-order interactions among the raw signal snapshots. We denote the feature set after the $\ell$-th layer as $\mathbf{S}^{(\ell)}\triangleq[\mathbf{s}_{1}^{(\ell)},\mathbf{s}_{2}^{(\ell)},\ldots,\mathbf{s}_{T}^{(\ell)}]$, which is also an unordered set. The self-attention block, $\mathbb{R}^{2M\times T}\to\mathbb{R}^{2M\times T}$, takes the feature set $\mathbf{S}^{(\ell)}\in\mathbb{R}^{2M\times T}$ as input and is defined in line 7 of Algorithm 1, where $\mathbf{W}_{Q}^{(\ell)},\mathbf{W}_{Q}^{(\ell)},\mathbf{W}_{Q}^{(\ell)}\in\mathbb{R}^{d_{\text{attn}}\times 2M}$ are the trainable parameter matrices for $\ell=1,2,\ldots,L$ with $d_{\text{attn}}$ being the width of attention activation. In the $\ell$-th layer, the query, key, and value are respectively denoted by $\mathbf{Q}^{(\ell)},\mathbf{K}^{(\ell)},\mathbf{V}^{(\ell)}\in\mathbb{R}^{d_{\text{attn}}\times T}$. In the self-attention block, each element of the feature set, i.e., $\mathbf{s}_{t}^{(\ell)},\forall t\in\{1,2,\ldots T\}$, attends to all other elements in $\mathbf{S}^{(\ell)}$, generating attention weights $\mathrm{Softmax}(\frac{(\mathbf{K}^{(\ell)})^{T}\mathbf{Q}^{(\ell)}}{\sqrt{d_{\text{attn}}}})$ that encode the pairwise statistical relationships of the set elements of $\mathbf{S}^{(\ell)}$. In addition, we also adopt layer normalization [46], i.e., $\mathrm{LayerNorm}(\cdot)$, and residual connection [47]. It is also followed by a feed-forward block in line 9, which are fully-connected layers to to enhance the representation capability, with $\mathrm{ReLU(\cdot)=\max(0,\cdot)}$ being the rectified linear unit activation function.
@@
The self-attention block effectively constructs pairwise cross terms of the feature set elements in $\mathbf{S}^{(\ell)}$. Through the stacking of $L$ self-attention layers, Snap-TF progressively constructs higher-order interactions. The first layer output encodes pairwise interactions, which, when processed by subsequent layers, enables the formation of terms involving up to four elements, and so forth. As a result, $L$ stacked attention layers can capture interactions of order up to $2^{L}$, while $L=\lceil\log_{2}(k+1)\rceil$ layers are enough to capture $k$-th order interactions. As a result, $L=3$ layers can provide sufficient representational capacity for DOA estimation under non-Gaussian modulated symbols used in communications, which usually have no higher than eighth-order statistics [44]. This will be verified by the ablation studies in Section V-F, suggesting that very few layers of Snap-TF can already offer competitive performance.
@@ IV-B3 Global Mean Pooling and Output Block
After $L$ layers, we obtain the final feature set of $T$ elements $\mathbf{S}^{(L)}=[\mathbf{s}_{1}^{(L)},\mathbf{s}_{2}^{(L)},\ldots,\mathbf{s}_{T}^{(L)}]$. To produce DOA estimates, we aggregate these $T$ features into a fixed-size vector using global mean pooling, i.e., computing the mean across all columns of $\mathbf{S}^{(L)}$. This pooled feature vector $\mathbf{q}$ is then fed into a small fully-connected network for final DOA estimation. Global mean pooling ensures that the Snap-TF DOA estimator $g_{\mathcal{W}}(\cdot)$ satisfies permutation invariance. The mean pooling is column-symmetric, so the order of the columns of $\mathbf{S}^{(L)}$ does not change the result of pooling. Together with the permutation equivariance property of self-attention blocks, overall, Snap-TF’s output is invariant to permutations of the snapshots, formally guaranteeing $g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}})=g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}}\bm{\Pi}_{T}),\forall\bm{\Pi}_{T}\in\mathcal{P}_{T}$, where $\mathcal{P}_{T}$ here denotes the set of all $T\times T$ permutation matrices. Global mean pooling also enables generalization to different snapshot counts $T$. The network can be trained on one range of $T$ values and applied to scenarios with different $T$ without retraining, since pooling can naturally aggregate different numbers of snapshots.
After global mean pooling, we apply fully-connected layers to the feature vector $\mathbf{q}$ to estimate the DOAs. To handle the varying number of sources in practical systems, we set the output dimension to the maximum allowed number of sources $K_{\max}$. The model outputs $K_{\max}$ angle estimates, and we select the first $K$ elements corresponding to the actual number of sources $K$, which enables the model to accommodate varying numbers of sources without requiring architectural changes.
@@ IV-C Computational Complexity
Unlike subspace methods, the proposed Snap-TF algorithm completely eliminates not only the computationally expensive covariance (or cumulant matrix) reconstruction step, but also the prohibitive eigen-decomposition and exhaustive spectrum search steps. It only needs to process the raw received signal snapshots of SLAs with dimension $M$. This complexity reduction is particularly significant for large-scale SLAs. By contrast, computational bottleneck of the proposed Snap-TF algorithm lies in computing attention matrices $\mathbf{K}^{(\ell)T}\mathbf{Q}^{(\ell)}$ with complexity $O(T^{2}d_{\text{attn}})$, where $d_{\text{attn}}$ is a hyper-parameter often set slightly larger than $N$. The complexity of Snap-TF algorithm is roughly linear with respect to $N$. Also, the number of snapshots $T$ in practical communication systems is usually small due to the short coherence interval. In addition, techniques to accelerate the attention step is extensively studied in the literature, e.g., linear attention [48], which scales linearly with $T$ and can further reduce the complexity. More importantly, Snap-TF enables the algorithm to capture the $k$-th order statistics using only $L=\text{$\lceil\log_{2}(k+1)\rceil$}$ layers.Hence, the overall complexity is roughly $O(\lceil\log_{2}(k+1)\rceil T^{2}d_{\text{attn}})$, which is significantly smaller compared to existing DL-based subspace methods. The advantage is particularly large for large-scale arrays and when higher-order statistics are utilized.
@@ V-A1 System Settings
We consider the uplink of a cellular wireless system where a multi-antenna BS attempts to localize multiple sources based on the unknown uplink data symbols, where sensing happens at the same time with data transmission. Specifically, we first consider a multi-user MIMO system where the BS employs an SLA with $M=5$ physical antennas and index set $\bm{\Omega}=\{1,2,5,8,10\}$, which is a 5-element MRA [13] with an aperture of $N=10$ half-wavelength units. Its difference co-array will lead to a 10-element virtual ULA that can localize $K_{\max}=N-1=9$ sources using subspace methods. In addition, to illustrate the scalability of the proposed Snap-TF algorithm, we also perform simulations for a large-scale sparse MIMO array which is configured as a 15-element MRA with $M=15$, $N=79$, and index set $\bm{\Omega}=\{1,2,3,6,11,16,27,38,49,60,66,72,78,79\}$, corresponding to a 79-element virtual ULA [43]. The number of signal snapshots is set as $T=50$, and the number of radio frequency (RF) chains is $N_{\text{RF}}=M$, if not otherwise specified. We will also study the generalization performance to both a smaller and a larger number of snapshots later. We consider the general case of random received signal power. The path losses $p_{k}$ are generated randomly and satisfy $\frac{\max_{k}|p_{k}|}{\min_{k}|p_{k}|}\leq 10$. The average received signal-to-noise-ratio (SNR) is $\text{SNR}\triangleq 10\log_{10}(\frac{\frac{1}{K}\sum_{k=1}^{K}|p_{k}|}{\eta})$ dB. During the training stage, we generate a mixture of received signals of different SNR levels that are sampled uniformly from a range of $-20$ to $20$ dB. For the Gaussian signal case, the signals are generated with $s_{k}(t)\sim\mathcal{CN}(0,1)$, while for the non-Gaussian data symbols, we mainly consider 16QAM symbols, which are drawn randomly from alphabet sets $\mathcal{S}_{\text{16QAM}}=\{a+jb\mid a,b\in\{\pm\frac{1}{\sqrt{10}},\pm\frac{3}{\sqrt{10}}\}\}$. For any value of $K$, each source DOA is sampled using $\theta_{k}\sim\mathcal{U}(\theta_{\min},\theta_{\max}),\,\forall k\in\{1,2,\ldots,K\}$, where we set $\theta_{\min}=\frac{\pi}{6}$, $\theta_{\max}=\frac{5\pi}{6}$, and the minimum angle separation as $\Delta_{\bm{\theta},\text{min}}=\frac{\pi}{60}$, respectively, if not otherwise specified. This relatively small minimum angle separation corresponds to a practical and challenging scenario. The wavelength is set as $\lambda=0.02$ meter, corresponding to 15 GHz carrier frequency.
@@ V-A2 Neural Network Settings
For the benchmark algorithms (DCR-T [30], DCR-G-Fro [29], DCR-G-Aff [29], and SRL-GD [31]), we employ the same wide residual network (WRN) as introduced in [31] with 16 layers and a widening factor of 8, i.e., the WRN-16-8 model, implemented by the authors’ open-source toolbox. All benchmarks adopt the same backbone architecture. Their differences mainly lie in the loss function and the output layer, which is an affine mapping with its output dimension customized to each approach, as introduced in Section III. The input to the network is the real-valued representation of the sample SCM of the SLA $\tilde{\mathbf{R}}_{\bm{\Omega}}\in\mathbb{C}^{M\times M}$, obtained by concatenating the real and imaginary parts, with a shape of $\mathbb{R}^{2\times M\times M}$. The output of the neural network is the real-valued representation of the reconstructed SCM of the virtual ULA $\hat{\mathbf{R}}\in\mathbb{C}^{N\times N}$, with a shape of $\mathbb{R}^{2\times N\times N}$. Same as [31], the reconstructed SCM is then fed into the Root-MUSIC algorithm to obtain the estimated DOAs.
@@
For the proposed Snap-TF, we set the attention activation dimension as $d_{\text{attn}}=96$, and adopt $L=2$ transformer layers for the Gaussian signals and $L=3$ layers in the case of non-Gaussian modulated symbols. After the transformer layers, we apply an affine mapping with a hidden dimension of 200 to transform the features to the estimated DOAs.

### v2

@@ IV-B Architecture of the Proposed Snap-TF
Next, we introduce the proposed Snap-TF architecture for multi-source DOA estimation. Its pseudo-codes are shown in Algorithm 1, while its schematic diagram is illustrated in Fig. 4. We discuss the details as follows.
Figure 4: Schematic diagram of the proposed Snap-TF algorithm.
@@ IV-B1 Signals as Tokens
We initialize by treating each raw snapshot as an input token to the transformer network. Specifically, for an $M$-element SLA, each snapshot $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$ is split into its real and imaginary parts and concatenated into a real-valued $2M$-dimensional vector. Stacking $T$ such vectors forms the initial feature set $\mathbf{S}^{(0)}\in\mathbb{R}^{2M\times T}$, whose $t$-th column corresponds to the $t$-th snapshot. We bypass the tokenization procedures in vanilla transformers [51] to let the network directly process higher-order interactions between raw signal snapshots to learn the higher-order statistics that are important for DOA estimation under non-Gaussian modulated symbols. Also, we deliberately drop the positional encoding and masking commonly employed in vanilla transformers [51] to preserve the permutation invariance of the architecture with respect to the ordering of signal snapshots $\mathbf{y_{\bm{\Omega}}}(t)$. Since a permutation of the snapshot index should not affect the output, we treat $\mathbf{S}^{(0)}$ as the initial feature set whose columns are the unordered set elements.
@@ IV-B2 Self-Attention Blocks for Higher-Order Interactions
Modulated data symbols carry higher-order statistics that can enhance localization capability, yet classical subspace methods require constructing and decomposing large and potentially redundant cumulant matrices. Moreover, the arrangement of higher-order terms directly determines the achievable localization capability [52]. In practice, with a limited number of snapshots, the optimal arrangement depends on many factors, calling for adaptive and data-driven computation. Higher-order cumulants are permutation-invariant symmetric polynomials of the set elements of $\mathbf{S}^{(0)}$. Since self-attention can effectively approximate such polynomials with theoretical guarantees [53], stacked self-attention layers in Snap-TF can implicitly and adaptively approximate the necessary higher-order information without explicitly reconstructing cumulant matrices, yielding compact features for the downstream DOA estimator.
@@
Snap-TF employs a stack of single-head self-attention blocks to capture the higher-order interactions among the raw signal snapshots. We denote the feature set after the $\ell$-th layer as $\mathbf{S}^{(\ell)}\triangleq[\mathbf{s}_{1}^{(\ell)},\mathbf{s}_{2}^{(\ell)},\ldots,\mathbf{s}_{T}^{(\ell)}]$, which is also an unordered set. The self-attention block, $\mathbb{R}^{2M\times T}\to\mathbb{R}^{2M\times T}$, takes the feature set $\mathbf{S}^{(\ell)}\in\mathbb{R}^{2M\times T}$ as input and is defined in line 7 of Algorithm 1, where $\mathbf{W}_{Q}^{(\ell)},\mathbf{W}_{K}^{(\ell)},\mathbf{W}_{V}^{(\ell)}$ are the trainable parameter matrices for $\ell=1,2,\ldots,L$ with $d_{\text{attn}}$ being the width of attention activation. In the $\ell$-th layer, the query, key, and value are respectively denoted by $\mathbf{Q}^{(\ell)},\mathbf{K}^{(\ell)},\mathbf{V}^{(\ell)}\in\mathbb{R}^{d_{\text{attn}}\times T}$. In the self-attention block, each element of the feature set, i.e., $\mathbf{s}_{t}^{(\ell)},\forall t\in\{1,2,\ldots T\}$, attends to all other elements in $\mathbf{S}^{(\ell)}$, generating attention weights $\mathrm{Softmax}(\frac{(\mathbf{K}^{(\ell)})^{\mathsf{T}}\mathbf{Q}^{(\ell)}}{\sqrt{d_{\text{attn}}}})$ that encode the pairwise statistical relationships of the set elements of $\mathbf{S}^{(\ell)}$. For example, define $\mathbf{q}_{i}^{(\ell)}=\mathbf{W}_{Q}^{(\ell)}\mathbf{s}_{i}^{(\ell)}$ and $\mathbf{k}_{j}^{(\ell)}=\mathbf{W}_{K}^{(\ell)}\mathbf{s}_{j}^{(\ell)}$. Their attention weight is $\mathrm{softmax}((\mathbf{k}_{j}^{(\ell)})^{\mathsf{T}}\mathbf{q}_{i}^{(\ell)}/\sqrt{d_{\mathrm{attn}}})$ which depends on the two set elements $\mathbf{s}_{i}^{(\ell)}$ and $\mathbf{s}_{j}^{(\ell)}$, and is thus adaptive to the input of the model. Self-attention hence creates adaptively weighted higher-order interactions without explicitly constructing cumulant matrices. In addition, we also adopt layer normalization [54], i.e., $\mathrm{LayerNorm}(\cdot)$, and residual connection [55]. It is also followed by a feed-forward block in line 9, which are fully-connected (FC) layers to to enhance the representation capability, with $\mathrm{ReLU(\cdot)=\max(0,\cdot)}$ being the rectified linear unit activation function.
@@
Through the stacking of $L$ self-attention layers, Snap-TF progressively constructs higher-order interactions. The first layer output encodes pairwise interactions, which, when processed by subsequent layers, enables the formation of terms involving up to four elements, and so forth. In addition, as seen from Fig. 4, while the self-attention operation creates higher-order interactions, the residual link preserves the lower-order interactions computed in previous layer. As a result, $L$ stacked attention layers can capture interactions of order up to $2^{L}$ and preserves all lower-order information, so $L=\lceil\log_{2}(k+1)\rceil$ layers are enough to capture $k$-th order interactions. For example, $L=3$ layers can provide sufficient representational capacity for DOA estimation under non-Gaussian modulated symbols used in communications, whose non-Gaussianity is well characterized by cumulants up to the eighth order [50]. This will be verified by the ablation studies in Section V-F, suggesting that very few layers of Snap-TF can already offer competitive performance.
@@ IV-B3 Global Mean Pooling and Output Block
After $L$ layers, we obtain the final feature set of $T$ elements $\mathbf{S}^{(L)}=[\mathbf{s}_{1}^{(L)},\mathbf{s}_{2}^{(L)},\ldots,\mathbf{s}_{T}^{(L)}]$. To produce DOA estimates, we aggregate these $T$ features into a fixed-size vector using global mean pooling, i.e., computing the mean across all columns of $\mathbf{S}^{(L)}$. This pooled feature vector $\mathbf{q}$ is then fed into a small FC network for final DOA estimation. Global mean pooling ensures that the Snap-TF DOA estimator $g_{\mathcal{W}}(\cdot)$ satisfies permutation invariance. The mean pooling is column-symmetric, so the order of the columns of $\mathbf{S}^{(L)}$ does not change the result of pooling. Together with the permutation equivariance property of self-attention blocks, overall, Snap-TF’s output is invariant to permutations of the snapshots, formally guaranteeing $g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}})=g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}}\bm{\Pi}_{T}),\forall\bm{\Pi}_{T}\in\mathcal{P}_{T}$, where $\mathcal{P}_{T}$ here denotes the set of all $T\times T$ permutation matrices. Global mean pooling also enables generalization to different numbers of snapshots $T$. The network can be trained on one range of $T$ values and applied to scenarios with different $T$ without retraining, since pooling can naturally aggregate different numbers of snapshots.
After global mean pooling, we apply FC layers to the feature vector $\mathbf{q}$ to estimate the DOAs. To handle the varying number of sources [^5] In 3GPP NR uplink, the base station either schedules uplink data on the physical uplink shared channel (PUSCH) by sending downlink control information (DCI) on the physical downlink control channel (PDCCH) addressed to a specific user via a radio network temporary identifier (RNTI), or enables configured-grant resources by radio resource control [56, 57, 58, 59]. Hence, the number of sources $K$ is available at the BS. $K$ in practical systems, we set the output dimension to the maximum allowed number of sources $K_{\max}$. The model outputs $K_{\max}$ angle estimates, and we select the first $K$ elements corresponding to the actual number of sources $K$, which enables the model to accommodate varying numbers of sources without requiring architectural changes.
@@ IV-C Computational Complexity
Unlike subspace methods, the proposed Snap-TF algorithm completely eliminates not only the computationally expensive covariance (or cumulant matrix) reconstruction step, but also the prohibitive eigen-decomposition and exhaustive spectrum search steps. It only needs to process the raw received signal snapshots of SLAs with dimension $M$. This complexity reduction is particularly significant for large-scale SLAs. By contrast, computational bottleneck of the proposed Snap-TF algorithm lies in computing attention matrices $\mathbf{K}^{(\ell)\mathsf{T}}\mathbf{Q}^{(\ell)}$ with complexity $O(T^{2}d_{\text{attn}})$, where $d_{\text{attn}}$ is a hyper-parameter often set slightly larger than $N$. The complexity of Snap-TF algorithm is roughly linear with respect to $N$. Also, the number of snapshots $T$ in practical communication systems is usually small due to the short coherence interval. In addition, techniques to accelerate the attention step is extensively studied in the literature, e.g., linear attention [61], which scales linearly with $T$ and can further reduce the complexity. More importantly, Snap-TF enables the algorithm to capture the $k$-th order statistics using $L=\text{$\lceil\log_{2}(k+1)\rceil$}$ layers. Hence, the overall complexity is roughly $O(\lceil\log_{2}(k+1)\rceil T^{2}d_{\text{attn}})$, typically smaller than existing DL-based subspace methods introduced in Section III. The advantage is particularly large for large-scale arrays, small number of snapshots $T$, and when higher-order statistics are used.
@@ V-A1 System Settings
We consider the “sensing for free” setup as introduced in Remark 1. Specifically, we first consider a multi-user MIMO system where the BS employs an SLA with $M=5$ physical antennas and index set $\bm{\Omega}=\{1,2,5,8,10\}$, which is a 5-element MRA [^6] We use MRAs to illustrate simulation results without loss of generality. The studied algorithms are geometry-agnostic as they exploit the subspace structure of received signal statistics, which holds for any array geometry. Different geometries yield different effective degree of freedom through their unique lags, but the algorithmic principles remain broadly applicable. [14] with an aperture of $N=10$ half-wavelength units. Its difference co-array will lead to a 10-element virtual ULA that can localize $K_{\max}=N-1=9$ sources using subspace methods. In addition, to illustrate the scalability of the proposed Snap-TF algorithm, we also perform simulations for a large-scale sparse MIMO array which is configured as a 15-element MRA with $M=15$, $N=79$, and index set $\bm{\Omega}=\{1,2,3,6,11,16,27,38,49,60,66,72,78,79\}$, corresponding to a 79-element virtual ULA [60]. The number of signal snapshots is set as $T=50$, and the number of radio frequency (RF) chains is $N_{\text{RF}}=M$, if not otherwise specified. We will also study the generalization performance to both a smaller and a larger number of snapshots later. We consider the general case of random received signal power. For each realization, the path losses are generated as i.i.d. preliminary samples $\tilde{p}_{k}\sim\mathcal{U}(p_{\min},p_{\max})$ with $p_{\max}/p_{\min}\leq 10$, and then rescaled to enforce unit mean power per source via $p_{k}\leftarrow K\,\tilde{p}_{k}/\sum_{j=1}^{K}\tilde{p}_{j}$, which ensures $\frac{\max_{k}p_{k}}{\min_{k}p_{k}}\leq 10$. The average received signal-to-noise ratio (SNR) is $\text{SNR}\triangleq 10\log_{10}\!\big(\frac{\frac{1}{K}\sum_{j=1}^{K}p_{k}}{\eta}\big)$ dB. During the training stage, we generate a mixture of received signals with SNR levels sampled uniformly from [20,20][-20,20] dB and set $\eta$ accordingly to match the target SNR in each realization. For the Gaussian signal case, the signals are generated with $s_{k}(t)\sim\mathcal{CN}(0,1)$, while for the non-Gaussian data symbols, we mainly consider 16QAM symbols, which are drawn randomly from alphabet sets $\mathcal{S}_{\text{16QAM}}=\{a+jb\mid a,b\in\{\pm\frac{1}{\sqrt{10}},\pm\frac{3}{\sqrt{10}}\}\}$. The sources signals are mutually independent and also uncorrelated with the noise if not otherwise stated, except for Section V-G where we study the special case of coherence sources. For a given $K$, we generate $\{\theta_{k}\}_{k=1}^{K}$ uniformly over $\theta_{\min},\theta_{\max}]$ with a minimum angle separation constraint $\Delta_{\bm{\theta},\text{min}}\triangleq\min_{i\neq j}\left|\theta_{i}-\theta_{j}\right|,\forall i,j\in\{1,2,\ldots,K\}$. This is realized by using the Poisson disk sampling [62], which yields an approximately uniform set over the interval while enforcing a minimum angle separation constraint $\Delta_{\bm{\theta},\text{min}}$. In the simulations, we set $\theta_{\min}=\tfrac{\pi}{6}$, $\theta_{\max}=\tfrac{5\pi}{6}$, and $\Delta_{\bm{\theta},\text{min}}=\tfrac{\pi}{60}$, if not otherwise stated. This relatively small minimum angle separation corresponds to a practical and challenging scenario. The wavelength is set as $\lambda=0.02$ meter, corresponding to 15 GHz carrier frequency in the upper mid-band spectrum [63].
@@ V-A2 Neural Network Settings
For the benchmark algorithms (DCR-T [33], DCR-G-Fro [32], DCR-G-Aff [32], and SRL-GD [34]), we employ the same wide residual network (WRN) as introduced in [34] with 16 layers and a widening factor of 8, i.e., the WRN-16-8 model, implemented by the authors’ open-source toolbox. All benchmarks adopt the same backbone architecture. Their differences mainly lie in the loss function and the output layer, which is an affine mapping with its output dimension customized to each approach, as introduced in Section III. The input to the network is the real-valued representation of the sample SCM of the SLA $\tilde{\mathbf{R}}_{\bm{\Omega}}\in\mathbb{C}^{M\times M}$, obtained by concatenating the real and imaginary parts, with a shape of $\mathbb{R}^{2\times M\times M}$. The output of the neural network is the real-valued representation of the reconstructed SCM of the virtual ULA $\hat{\mathbf{R}}\in\mathbb{C}^{N\times N}$, with a shape of $\mathbb{R}^{2\times N\times N}$. Same as [34], the reconstructed SCM and the number of sources $K$ are then fed into the Root-MUSIC algorithm to obtain the estimated DOAs. All the benchmarks, as well as our proposed algorithm, use the number of sources $K$ as an input, which is available at the BS according to 3GPP standards, as explained before in footnote 5.
@@
For the proposed Snap-TF, we set the attention activation dimension as $d_{\text{attn}}=96$, and adopt $L=2$ transformer layers for the Gaussian signals and $L=3$ layers in the case of non-Gaussian modulated symbols. After the transformer layers, we apply an affine mapping with a hidden dimension of 200 to transform the features to the estimated DOAs. Specifically, for the proposed Snap-TF method, we generate the DOA label by first sorting the $K$ angles in ascending order, and then zero-pad the rest $(K_{\text{max}}-K)$ elements. For example, let $K_{\text{max}}=9$ and $K=3$ with actual DOAs $\theta_{1}=\frac{\pi}{6}$, $\theta_{2}=\frac{\pi}{3}$, and $\theta_{3}=\frac{\pi}{2}$, the DOA label will be generated as $\frac{\pi}{6},\frac{\pi}{3},\frac{\pi}{2},0,0,0,0,0,0]$.